OpenStudy (anonymous):

Using complete sentences, compare and contrast the graphs of y = x2 − 36 and y + x2 = 8x − 12.

OpenStudy (precal):

y=x^2-36 is down 36 units y=-x^2+8x-12 reflected across x axis I think it is right 4 units and up 4 units